OpenStudy (anonymous):

Given the function f(x) = 2(x + 6), find x if f(x) = 22.

OpenStudy (saifoo.khan):

f(x) = 2(x + 6) 22 = 2(x+6) Solve. :)

OpenStudy (saifoo.khan):

nope!

jhonyy9 (jhonyy9):

candyme not x=22 looke it again f(x)=22

OpenStudy (anonymous):

2(x+6)=22 Make the x a 22...So it will now be 2(22+6)=22 Distribute next by multiplying  Wanted to show the arrows.. Hope this helped!